On August 1, Rantoul Stores Inc. is considering leasing a building and purchasing the necessary equipment to operate a retail store. Alternatively, the company could use the funds to invest in $162,000 of 5% U.S. Treasury bonds that mature in 16 years. The bonds could be purchased at face value. The following data have been assembled:

Cost of store equipment $162,000
Life of store equipment 16 years
Estimated residual value of store equipment $12,800
Yearly costs to operate the store, excluding depreciation of store equipment $61,795
Yearly expected revenues—years 1–8 $83,600
Yearly expected revenues—years 9–16 $74,700
Required:
1. Prepare a differential analysis as of August 1 presenting the proposed operation of the store for the 16 years (Alternative 1) as compared with investing in U.S. Treasury bonds (Alternative 2). Refer to the lists of Labels and Amount Descriptions for the exact wording of the answer choices for text entries. For those boxes in which you must enter subtracted or negative numbers use a minus sign. If there is no amount or an amount is zero, enter "0". A colon (:) will automatically appear if required.
2. Based on the results disclosed by the differential analysis, should the proposal be accepted?
3. If the proposal is accepted, what would be the total estimated income from operations of the store for the 16 years?
Answer #1
DIFFERENTIAL ANALYSIS IN $
ALTERNATIVE 1 ALTERNATIVE 2 DIFFERENTIAL EFFECT OF ALT 1 AND ALT 2
REVENUE [NOTE 1] 1266400

129600

(1136800)

LESS:

OPERATING COST OF STORE [NOTE 2] 988720 0 988720
COST OF EQUIPMENT LESS SALVAGE VALUE [NOTE 3] 149200 0 149200
PROFIT /(LOSS) 128480 129600 (1120)

NOTE 1

ALT 1

[83600*8] + [74700*8] = 668800+597600

=1266400

ALT 2

162000*5%*16YEARS

= 129600

COST OF OPERATING STORE NOTE 2

61795*16YEARS =988720

NOTE 3

COST OF EQUIPMENT WHICH IS IRRECOVERABLE

=162000-12800

=149200$

2. NO PROPOSAL SHOULD NOT BE ACCEPTED AS IT WOULD RESULT IN 1120$ LOSS AS COMPARED TO WHEN INVESTED IN BONDS.

3.ESTIMATED INCOME

IN $
ESTIMATED REVENUE 1266400
LESS
OPERATING COST EXCLUDING DEPRECIATION 988720
COST OF EQUIPMENT LESS RESIDUAL VALUE 149200
ESTIMATED INCOME FROM STORE 128480